### v1

# A note for Carleson measure on bounded $\mathbb{C}$}-doconvex domains

Abstract.
Let $0<q<p<\infty$, $\Omega$ be a bounded $\mathbb{C}$}-convex domains in ${\mathbb{C}}^{n}$. We establish several equivalent characterizations for the boundedness of Carleson embedding $J_{\mu}:A_{\alpha}^{p}\hookrightarrow L^{q}(\mu)$ on $\Omega$ with sharp $\mathcal{B}$} condition. Furthermore, we prove that the boundedness of $J_{\mu}$ is equivalent to its compactness.
Key words and phrases:
Carleson measure, Weakly pseudoconvex domains, Sharp $\mathcal{B}$}, Bergman spaces
2020 Mathematics Subject Classification:
Primary 46E30; Secondary 32A25, 32A36
*:Corresponding author.
1. Introduction and main result
The notion of Carleson measure is an important object in complex analysis and function theory, which was introduced by Carleson [6] in his celebrated work on the corona problem in 1962. Specifically, Carleson characterized the boundedness of the Carleson embedding $J_{\mu}:H^{p}\hookrightarrow L^{p}(\mu)$ by using Carleson square $S_{\theta_{0},h}=\{re^{i\theta}\in{\mathbb{D}}:1-h\leq r<1,|\theta-\theta_{0}|\leq h\}$. Later, Duren [9] extended Carleson’s result to the boundedness of $J_{\mu}:H^{p}\hookrightarrow L^{q}(\mu)$ for $0<p\leq q<\infty$ by adapting the original argument, moreover, the Carleson square can be replaced by pseudohyperbolic disks. Luecking [15] obtained the boundedness of $J_{\mu}:H^{p}\to L^{q}(\mu)$ for the case of $p>q$ by using Khinchine’s inequality and the theory of tent spaces associated with Hardy spaces.
The extension of $q$-Carleson measure to Bergman spaces has attracted considerable attention from many scholars, it was characterized by Hastings [17] and Oleinik [18] for $p\leq q$, Luecking [16] for $p>q$. In higher dimension, Cima-Wogen [7] and Luecking-Duren-Weir [14, 10] stuied the case of unit ball ${\mathbb{B}}^{n}$ by using the Euclidean geometry of ${\mathbb{B}}^{n}$ and conformal invariants (such as Bergman or Kobayashi balls) repectively. In order to extend the above results to other domains of ${\mathbb{C}}^{n}$, Cima-Mercer and Abate-Raissy-Saracco characterrized the $q$-Carleson measure for $A^{p}$ on strongly pseudoconvex domains by using the Euclidean geometric quantities and intrinsic Kobayashi geometry respectively, see [1, 2, 3, 8] for the details. For the case of weakly pseudoconvex domains, Li-Liu-Wang [13] characterized the $2$-Carleson measurable for $A^{2}$ by using Kobayashi balls on bounded convex domains with smooth boundary of finite type. Bi-Su-Zhang [5] characterized the $2$-Carleson measure for $A^{2}$ and $A^{p}$ respectively on bounded $\mathbb{C}$}-convex domains.
Khanh-Liu-Thuc [11] introduced the sharp $\mathcal{B}$} type condition (see Sect.2.4 ) for Bergman kerenl. Roughly speaking, Bergman kernel satisfies the sharp $\mathcal{B}$} condition ensures that the Bergman projection is bounded on $L^{p}$ for $1<p<\infty$. More recently, there is a series of work involved function spaces and operator theory related such domains. For example, Khanh-Liu-Thuc-Tien [11, 12] characterrized the boundedness and Schatten class of Bergman-Toeplitz operators for weighted spaces on weakly pseudoconvex domains with Bergman kernel satisfies sharp $\mathcal{B}$} condition. Zhang [19] characterized the $q$-Carleson measure for the weighted Bergman spaces $A_{\alpha}^{p}$ on bounded $\mathbb{C}$}-convex domains with sharp $\mathcal{B}$} type condition, and what’s more, Zhang’s results extended the results in [5] to the case of $p\leq q$ and the weighted Bergman spaces $A_{\alpha}^{p}$, but for the case of $p>q$, Zhang didn’t characterized the full Carleson measure, Zhang’s result is stated as follows.
Theorem 1.1.
[19, Theorem 1.8] Let $\Omega\subset{\mathbb{C}}^{n}$ be a bounded $\mathbb{C}$}-convex domain whose Bergman kernel of sharp $\mathcal{B}$} type, $\mu$ be a finite positive Borel measure on $\Omega$. Assume that $0<q<p<\infty$, if $\mu$ satisfies one of the following three conditions
  • (i)
    $p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widetilde{\mu}_{t}(z)\in L^{\frac{p}{p-q}}$ for some $t>1$;
  • (ii)
    $\{p_{\Omega}(a_{j})^{2+\frac{2(\alpha-1)q}{p}}\widehat{\mu}_{r}(a_{j})\}_{j=1}^{\infty}\in l^{\frac{p}{p-q}}$ for some $r\in(0,1)$ and $r$-lattice $\{a_{j}\}_{j=1}^{\infty}$;
  • (iii)
    $p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widehat{\mu}_{R}\in L^{\frac{p}{p-q}}$ for some $R\in(0,1)$.
Then $J_{\mu}:A_{\alpha}^{p}\hookrightarrow L^{q}(\mu)$ is bounded.
Motivated by Theorem 1.1, we give the following main result in this paper. Theorem 1.2 gives an equivalent characterization of $q$-Carleson measure for $A_{\alpha}^{p}$, which generalizes the Theorem 1.1. Moreover, we also show that $J_{\mu}$ is bounded if and only if $J_{\mu}$ is compact for the case of $p>q$.
Theorem 1.2.
Let $\Omega$ be a bounded $\mathbb{C}$}-convex domain whose Bergman kernel of sharp $\mathcal{B}$}-type. Let $\mu$ be a finite positive Borel measure, $0<q<p<\infty$. Then the following are equivalent.
  • (i)
    $J_{\mu}:A_{\alpha}^{p}\hookrightarrow L^{q}(\mu)$ is compact;
  • (ii)
    $J_{\mu}:A_{\alpha}^{p}\hookrightarrow L^{q}(\mu)$ is bounded;
  • (iii)
    For some $r\in(0,1)$ and $r$-lattice $\{a_{j}\}_{j=1}^{\infty}$, we have
    $\left\{p_{\Omega}(a_{j})^{2+\frac{2q(\alpha-1)}{p}}\widehat{\mu}_{r}(a_{j})\right\}_{j=1}^{\infty}\in l^{\frac{p}{p-q}};$
  • (iv)
    $p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widehat{\mu}_{R}(z)\in L^{\frac{p}{p-q}}(\Omega)$ for some $R\in(0,1)$.
  • (v)
    $p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widehat{\mu}_{R}(z)\in L^{\frac{p}{p-q}}(\Omega)$ for all $R\in(0,1)$.
This paper is organized as follows. Some preliminaries are given in Section 2; An atomic type decomposition and the proof of Theorem 1.2 are showed in Section 3.
2. Preliminaries
2.1. Notations
  • For a domain $\Omega$, ${\mathcal{H}}(\Omega)$ denotes the set of holomorphic functions on $\Omega$.
  • $|\cdot|$ denote the standard Euclidean norm.
  • For $z\in\Omega\subsetneq{\mathbb{C}}^{n}$, $\delta_{\Omega}(z):=\inf\limits_{z_{0}\in\partial\Omega}|z-z_{0}|$.
  • $a\lesssim b$ or $a\gtrsim b$ means that there exists a positive constant $C$ such that $a\leq Cb$ or $a\geq Cb$, $a\asymp b$ if $a\lesssim b$ and $a\gtrsim b$.
2.2. Invariants
Given a domain $\Omega\subset\mathbb{C}^{n}$ (with $n\geq 2$), the (infinitesimal) Kobayashi metric is the pseudo-Finsler metric defined by
$\kappa_{\Omega}(x;v):=\inf\bigl\{|\xi|:f\in{\mathcal{H}}(\mathbb{D},\Omega),\text{ with }f(0)=x,\;(df)_{0}(\xi)=v\bigr\}.$
For any curve $\sigma:[a,b]\to\Omega$, the Kobayashi length is
$l_{\kappa}(\sigma):=\int_{a}^{b}\kappa_{\Omega}\bigl(\sigma(t);\sigma^{\prime}(t)\bigr)\,dt,$
and the Kobayashi pseudo-distance on $\Omega$ is
$d_{\Omega}^{K}(x,y):=$ $\inf_{\sigma}\bigl\{l_{\kappa}(\sigma)\mid\sigma:[a,b]\to\Omega\text{ is any absolutely continuous curve with }$
$\sigma(a)=x\text{ and }\sigma(b)=y\bigr\}.$
Suppose $\Omega$ is a bounded domains in ${\mathbb{C}}^{n}$, $z\in\Omega$ and $r\in(0,1)$. Let $B_{\Omega}(z,r)$ be the Kobayashi balls of centre $z$ and radius $\frac{1}{2}\log\frac{1+r}{1-r}$, i.e.
$B_{\Omega}(z,r)=\bigl\{z\in X\mid\tanh d_{\Omega}^{K}(z,w)<r\bigr\}.$
Note that $\rho_{\Omega}=\tanh d_{\Omega}^{K}$ is still a distance on $X$ since $\tanh$ is a strictly convex function on $\mathbb{R}^{+}$.
For a bounded pseudoconvex domains $\Omega\subset{\mathbb{C}}^{n}$, Set $A^{2}$ be the Hilbert spaces consists of all $f\in{\mathcal{H}}(\Omega)$ such that $\|f\|_{A^{2}}^{2}=\int_{\Omega}|f(z)|^{2}dV(z)<\infty$, where $dV(z)$ is the ordinary Lebesgue measure. Riesz’s theorem yields that there exists a unique reproducing kernel function $K_{\Omega}(z,w)$ such that
$f(z)=\int_{\Omega}f(w)\overline{K_{\Omega}(w,z)}dV(w)=\int_{\Omega}f(w)K_{\Omega}(z,w)dV(w),$
$K_{\Omega}(z,w)$ is called the Bergman kernel. For $\alpha\leq 0$, the weighted Lebesgue space $L_{\alpha}^{p}$ consists of all measurable functions such that
$\|f\|_{L_{\alpha}^{p}}^{p}=\int_{\Omega}|f(z)|^{p}K_{\Omega}(z,z)^{\alpha}dV(z)<\infty,$
$A_{\alpha}^{p}:=L_{\alpha}^{p}\bigcap{\mathcal{H}}(\Omega)$.
Let $X$ be the space consists of holomorphic functions, for $0<p,q<\infty$ and a positive Borel measure $\mu$, we say that $\mu$ is a $q$-Carleson measure for $X$ if there exists a finite constant $C>0$ such that
$\|f\|_{L^{q}(\mu)}\leq C\|f\|_{X},\quad f\in X,$
where $\|f\|_{L^{q}(\mu)}=\left(\int_{\Omega}|f(z)|^{q}d\mu\right)^{\frac{1}{q}}<\infty.$ It is clear that $\mu$ is a $q$-Carleson measure for $X$ if and only if the Carleson embedding
$J_{\mu}:X\hookrightarrow L^{q}(\mu)$
is bounded. What’s more, $\mu$ is a vanishing Carleson measure if the embedding $J_{\mu}$ is compact in the sense that $\lim\limits_{j\to\infty}\|f_{j}\|_{L^{q}(\mu)}=0$ whenever $\{f_{j}\}_{j=1}^{\infty}$ is a bounded sequence in $X$ which converges to 0 uniformly on any compact subset of $\Omega$. For any given $r\in(0,1)$, the averaging function of $\mu$ related to Kobayashi ball is defined as
$\widehat{\mu}_{r}(z):=\frac{\mu(B_{\Omega}(z,r))}{V(B_{\Omega}(z,r))},\quad z\in\Omega,$
where $V(B_{\Omega}(z,r))$ is the Euclidean volume of $B_{\Omega}(z,r)$.
2.3. Geometry of $\mathbb{C}$}-convex domains
Definition 2.1.
Suppose $\Omega$ is a domain in ${\mathbb{C}}^{n}$. For any affine complex lines $l$, if $l\bigcap\Omega$ is simply connected whenever $l\bigcap\Omega\neq\emptyset$, then $\Omega$ is said to a $\mathbb{C}$}-convex domain.
The following geometric objects related to an arbitrary domain $\Omega\subset\mathbb{C}^{n}$ containing no complex lines are needed in this paper. Let $q\in\Omega$, and let $\delta_{\Omega}(q)$ denote the distance from $q$ to the boundary $\partial\Omega$. Fix $q_{1}\in\partial\Omega$ so that
$\eta_{1}(q):=|q_{1}-q|=\delta_{\Omega}(q).$
Put
$H_{1}:=q+\operatorname{Span}(q_{1}-q)^{\perp}\quad\text{and}\quad\Omega_{1}:=\Omega\cap H_{1}.$
Let $q_{2}\in\partial\Omega_{1}$ so that
$\eta_{2}(q):=|q_{2}-q|=d_{\Omega_{1}}(q).$
Set
$H_{2}:=q+\operatorname{Span}(q_{1}-q,q_{2}-q)^{\perp},\quad\Omega_{2}:=\Omega\cap H_{2},$
and. Continuing this procedure we are led to an orthonormal basis
$e_{j}=\frac{q_{j}-q}{|q_{j}-q|},\quad 1\leq j\leq n,$
which is called minimal for $\Omega$ at $q$. It should be mentioned that after rotation we may always assume that $e_{1},e_{2},\dots,e_{n}$ is the standard basis of $\mathbb{C}^{n}$. Clearly, if $\Omega$ is a bounded domain, then for any $z\in\Omega$ we have
$\eta_{1}(z)\leq\eta_{2}(z)\leq\dots\leq\eta_{n}(z)\leq\operatorname{diam}(\Omega)<\infty,$
where $\operatorname{diam}(\Omega):=\sup\{|z-w|:z,w\in\Omega\}.$
The following lemma describes the geometric characteristics of Kobayashi balls on bounded $\mathbb{C}$}-convex domains.
Lemma 2.2.
[NIkolv2011, Theorem 1] Let $\Omega\subset\mathbb{C}^{n}$ be a $\mathbb{C}$-convex domain containing no complex lines, and $z\in\Omega$. Then for any $r\in(0,1)$,
$\mathbb{D}^{n}(z,c_{1}\eta(z))\subset B_{\Omega}(z,r)\subset\mathbb{D}^{n}(z,c_{2}\eta(z)),$
where $c_{i}$ $(i=1,2)$ is some constant depending only on $n$ and $r$, $\eta(z):=(\eta_{1}(z),\dots,\eta_{n}(z))$, and
$\mathbb{D}^{n}(z,\eta(z)):=\{w\in\mathbb{C}^{n}:|w_{i}-z_{i}|<\eta_{i}(z),\ i=1,\dots,n\}.$
For plurisubharmonic functions, the sub mean value inequality are valid for Kobayashi balls on bounded $\mathbb{C}$}-convex domains.
Lemma 2.3.
[5, Proposition 2.3] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded $\mathbb{C}$-convex domain. For any $r\in(0,1)$, we denote $R:=\frac{1}{2}(1+r)\in(0,1)$. Then for any $z_{0}\in\Omega$ and $z\in B_{\Omega}(z_{0},r)$, we have
$\varphi(z)\lesssim\frac{1}{V(B_{\Omega}(z_{0},r))}\int_{B_{\Omega}(z_{0},R)}\varphi dV$
for all nonnegative plurisubharmonic functions $\varphi$ on $\Omega$.
The following cover type lemma plays an significant role in the proof of Theorem 1.2.
Lemma 2.4.
[5, Lemma 2.4] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded $\mathbb{C}$-convex domain. For any $r\in(0,1)$, there exists a positive integer $m$ and a sequence $\{a_{j}\}_{j=0}^{\infty}\subset\Omega$ satisfy the following two conditions:
  • (i)
    $\Omega=\bigcup_{j=0}^{\infty}B_{\Omega}(a_{j},r)$;
  • (ii)
    Every point in $\Omega$ belongs to at most $m$ of the balls $B_{\Omega}(a_{j},R)$, where $R:=\frac{1}{2}(1+r)$.
Such a sequence $\{a_{j}\}_{j=0}^{\infty}\subset\Omega$ is said to be an $r$-lattice in $\Omega$.
The lower estimate for the Bergman kernel is given as follows.
Lemma 2.5.
[19, Lemma 2.4] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded $\mathbb{C}$-convex domain. Then for any $r\in(0,1)$, $z\in\Omega$ and $w\in B_{\Omega}(z,r)$, we have that
$|K_{\Omega}(z,w)|\gtrsim K_{\Omega}(z,z)^{\frac{1}{2}}K_{\Omega}(w,w)^{\frac{1}{2}},$
where the constant in $\gtrsim$ depends only on $r$ and the domain $\Omega$.
For $z\in\Omega$ and $r\in(0,1)$. Let $p_{\Omega}(z)=\prod\limits_{i=1}^{n}\eta_{i}(z)$, then by [19], we have
(2.1) $\frac{1}{C}p_{\Omega}(z)\leq p_{\Omega}(w)\leq Cp_{\Omega}(z),\quad w\in B_{\Omega}(z,r)$
for some positive constant $C=C(r)$.
By (2.1), the following estimate for Bergman kernel on diagonal is given.
Lemma 2.6.
[19, Lemma 2.5] Let $\Omega\subset\mathbb{C}^{n}$ be a $\mathbb{C}$-convex domain containing no complex lines. Then
$K_{\Omega}(z,z)\asymp\rho_{\Omega}(z)^{-2},\quad z\in\Omega.$
Remark.
By (2.1) and Lemma 2.6, for $w\in B_{\Omega}(z,r)$, we have
(2.2) $|K_{\Omega}(z,w)|\gtrsim K_{\Omega}(z,z).$
2.4. Bergman kernel with sharp $\mathcal{B}$} conditions
In this subsection, we introduce a sharp $\mathcal{B}$} condition for Bergman kerenl, we firstly recall that the definition of $\mathcal{B}$}- systems.
Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}$. For $z\in\overline{\Omega}$ near the boundary $\partial\Omega$, a family of functions $\mathcal{B}=\{b_{j}(z,\cdot)\}_{j=1}^{n}$ is called a $\mathcal{B}$-system at $z$ if there exist a neighbourhood $U$ of $z$ and a positive integer $m\geq 2$ such that for all $w\in U$,
$b_{1}(z,w):=\frac{1}{\delta(z,w)}\quad\text{and}\quad b_{j}(z,w):=\sum_{k=2}^{m}\left(\frac{A_{jk}(z)}{\delta(z,w)}\right)^{\frac{1}{k}},\quad\text{for }j=2,\ldots,n,$
where $A_{jk}$ are non-negative bounded functions, and $\delta(z,w)$ is the pseudo-distance between $z$ and $w$ defined as follows:
$\delta(z,w):=|\delta_{\Omega}(z)|+|\delta_{\Omega}(w)|+|z_{1}-w_{1}|+\sum_{l=2}^{n}\sum_{s=2}^{m}A_{ls}(z)|z_{l}-w_{l}|^{s}$
For $z\in\overline{\Omega}$ near the boundary, the Bergman kernel $K_{\Omega}$ is said to be of $\mathcal{B}$-type at $z\in\overline{\Omega}$ near the boundary $\partial\Omega$ if there exist positive constants $c$ and $C$ such that for all $w\in\Omega\cap\mathbb{B}(z,c)$,
$|K_{\Omega}(z,w)|\leq C\prod_{j=1}^{n}b_{j}^{2}(z,w).$
The Bergman kernel $K_{\Omega}$ is said to be of sharp $\mathcal{B}$-type at $z$ if $K_{\Omega}$ is of $\mathcal{B}$-type and has the sharp lower-bound on diagonal, i.e.
$K_{\Omega}(z,z)\asymp\prod_{j=1}^{n}b_{j}^{2}(z,z).$
We finish this section with a Forelli-Rudin estimate on domains with Bergman kernel satisfying the sharp $\mathcal{B}$} condition, which will be used to construct an atomic type decomposition.
Lemma 2.7.
[19, Proposition 2.7] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain whose Bergman kernel $K_{\Omega}$ is of sharp $\mathcal{B}$-type. Assume that $\alpha\leq 0,\,\beta\geq 0,\,t+\alpha\geq 1$ and $1<\beta-2\alpha<2t-2$. Then the following inequality
$\int_{\Omega}|K_{\Omega}(z,w)|^{t}K_{\Omega}(w,w)^{\alpha}d_{\Omega}(w)^{\beta}dV(w)\lesssim K_{\Omega}(z,z)^{t+\alpha-1}d_{\Omega}(z)^{\beta}$
holds for every $z\in\Omega$.
3. Proof of main result
In this section, we give the proof of Theorem 1.2. In order to prove theorem 1.2, we need to construct atomic type decomposition of $A_{\alpha}^{p}$, which is stated as follows.
Theorem 3.1.
Let $r\in(0,1)$ be fixed, and $\{w_{k}\}$ be an $r$-lattice. For $0<p<\infty$ and $\{c_{k}\}\in l^{p}$, $\alpha+Np>1$. Set $f(z)=\sum_{k=1}^{\infty}c_{k}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}K_{\Omega}(z,w_{k})^{N}$,
$\|f\|_{A_{\alpha}^{p}}\lesssim 1.$
The following lemma can be obtained by Lemma 2.7 and Schur’s test, it can be used to construction of atomic decomposition in the proof of Theorem 3.1.
Lemma 3.2.
Let $\Omega$ be a bounded pseudoconvex domain whose Bergman kernel $K_{\Omega}(z,w)$ satisfying sharp $\mathcal{B}$} condition. Assume that
(3.1) $\begin{cases}&q+p\sigma+\alpha\leq 0,\\ &s+q+p\sigma+\alpha\geq 1,\\ &-1<-2(q+p\sigma+\alpha)<2s-2,\\ &t+p^{\prime}\sigma\leq 0,\\ &s+t+p^{\prime}\sigma\geq 1,\\ &-1<-2(t+p^{\prime}\sigma+\alpha)<2s-2,\\ &q+s+t-1\leq 0.\end{cases}$
Then the integral operator
$f\mapsto T_{q,s,t}f(z)=K_{\Omega}(z,z)^{q}\int_{\Omega}K_{\Omega}(z,w)^{s}K_{\Omega}(w,w)^{t}f(w)dV(w)$
is bounded on $L^{p}(dV_{\alpha})$.
Proof.
To prove the boundedness of $T_{q,s,t}$, we only prove the boundedness $S_{q,s,t}f(z)=K_{\Omega}(z,z)^{q}\int_{\Omega}|K_{\Omega}(z,w)|^{s}K_{\Omega}(w,w)^{t}f(w)dV(w)$. To do this, let $h(z)=K_{\Omega}(z,z)^{\sigma}$,
$K_{\Omega}(z,z)^{q}\int_{\Omega}|K_{\Omega}(z,w)|^{s}K_{\Omega}(w,w)^{t}dV(w)$
$=K_{\Omega}(z,z)^{q}\int_{\Omega}|K_{\Omega}(z,w)|^{s}K_{\Omega}(w,w)^{t-\alpha}dV_{\alpha}(w).$
On the one hand, by Lemma 2.7 and
$K_{\Omega}(z,z)^{q}\int_{\Omega}|K_{\Omega}(z,w)|^{s}K_{\Omega}(w,w)^{t-\alpha}K_{\Omega}(w,w)^{p^{\prime}\sigma}dV_{\alpha}(w)$
$=K_{\Omega}(z,z)^{q}\int_{\Omega}|K_{\Omega}(z,w)|^{s}K_{\Omega}(w,w)^{t+p^{\prime}\sigma}dV(w)$
$\lesssim K_{\Omega}(z,z)^{q+s+t-1+p^{\prime}\sigma}$
$\leq CK(z,z)^{p^{\prime}\sigma}=Ch(z)^{p^{\prime}}.$
On the other hand,
$\int_{\Omega}K_{\Omega}(z,z)^{q}|K_{\Omega}(z,w)|^{s}K_{\Omega}(w,w)^{t-\alpha}K_{\Omega}(z,z)^{p\sigma}dV_{\alpha}(z)$
$=K_{\Omega}(w,w)^{t-\alpha}\int_{\Omega}|K_{\Omega}(z,w)|^{s}K_{\Omega}(z,z)^{q+\alpha+p\sigma}dV(z)$
$\lesssim K_{\Omega}(z,z)^{t+s+p\sigma-1}$
$\leq CK_{\Omega}(w,w)^{p\sigma}$
$=Ch(w)^{p}.$
Applying Schur’s test, we obtain that $S_{q,s,t}$ is bounded on $L^{p}(dV_{\alpha})$, the proof is completed. ∎
Now we give the proof of Theorem 3.1.
Proof of Theorem 3.1.
It is obvious that the $f$ is holomorphic. For $Np+\alpha>1$, by Lemma 2.7, we have
(3.2) $\|K_{\Omega}(z,w_{k})^{N}\|_{A_{\alpha}^{p}}\lesssim K_{\Omega}(w_{k},w_{k})^{\frac{\alpha-1}{p}+N}.$
For $0<p\leq 1$, using the fact that
$\left(\sum_{j=1}^{\infty}b_{j}\right)^{p}\leq\sum_{k=1}^{\infty}b_{j}^{p}\,\,\,\text{for}\,\,\,b_{j}\geq 0,$
(3.2) yields that
$\|f\|_{A_{\alpha}^{p}}^{p}\lesssim\sum_{k=0}^{\infty}|c_{k}|^{p}K_{\Omega}(w_{k},w_{k})^{1-\alpha-pN}\|K_{\Omega}(\cdot,w_{k})^{N}\|_{A_{\alpha}^{p}}^{p}\lesssim\sum_{k=1}^{\infty}|c_{k}|^{p}.$
For $1<p<\infty$, Set
$\begin{cases}&D_{k}=B_{\Omega}(w_{k},r),\\ &F(z)=\sum_{k=0}^{\infty}|c_{k}|v_{\alpha}(D_{k})^{-\frac{1}{p}}\chi_{k}(z),\end{cases}$
then we have
$\|F\|_{A_{\alpha}^{p}}^{p}\lesssim\|c_{k}\|_{l^{p}}^{p}.$
Let q=1st,s=N,t=1Nq=1-s-t,s=N,t=1-N in (3.1), then by lemma 2.2, lemma 2.5 and (2.1), we have
$S_{q,s,t}F(z)$
$=K_{\Omega}(z,z)^{1-N-(1-N)}\int_{\Omega}|K_{\Omega}(z,w)|^{N}K_{\Omega}(w,w)^{1-N}\left(\sum_{k=0}^{\infty}|c_{k}|v_{\alpha}(D_{k})^{-\frac{1}{p}}\chi_{k}(w)\right)dv(w)$
$=\sum_{k=0}^{\infty}|c_{k}|v_{\alpha}(D_{k})^{-\frac{1}{p}}\int_{D_{k}}|K_{\Omega}(z,w)|^{N}K_{\Omega}(w,w)^{N}dv(w)$
$\gtrsim\sum_{k=0}^{\infty}|c_{k}|K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}}\int_{D_{k}}|K_{\Omega}(z,w)|^{N}K_{\Omega}(w,w)^{1-N}dv(w)$
$\gtrsim\sum_{k=0}^{\infty}|c_{k}|K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}+1-N}\int_{D_{k}}|K_{\Omega}(z,w)|^{N}dv(w).$
By Lemma 2.3,
$|K_{z}(w_{k})|^{N}$ $\lesssim\frac{1}{V(D_{k})}\int_{D_{k}}|K_{\Omega}(z,w)|^{N}dv(w)$
$\lesssim p_{\Omega}(w_{k},w_{k})^{-2}\int_{D_{k}}|K_{\Omega}(z,w)|^{N}dv(w)$
$\asymp K_{\Omega}(w_{k},w_{k})\int_{D_{k}}|K_{\Omega}(z,w)|^{N}dv(w),$
then
$S_{q,s,t}F(z)\gtrsim\sum_{k=0}^{\infty}|c_{k}|K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}\gtrsim|f(z)|.$
Hence, lemma 3.1 that yields
$\|f\|_{\alpha}^{p}\lesssim\|S_{q,s,t}F\|_{L_{\alpha}^{p}}^{p}\lesssim\|F\|_{L_{\alpha}^{p}}^{p}\lesssim\|c_{k}\|_{l^{p}}^{p}.$
The proof is completed. ∎
In order to prove (iv) $\Longleftrightarrow$ (v) in Theorem LABEL:restate, we need the following results.
Lemma 3.3.
[19, Lemma 5.3] Let $1<p<\infty$ and $s\in\mathbb{R}$. Then the function $p_{\Omega}^{s}\widehat{\mu}_{r}\in L^{p}(\Omega)$ for some $r\in(0,1)$ if and only if $p_{\Omega}^{s}\widehat{\mu}_{R}\in L^{p}(\Omega)$ for all $R\in(0,1)$. Moreover, we have
$\|p_{\Omega}^{s}\widehat{\mu}_{r}\|_{p}\asymp\|p_{\Omega}^{s}\widehat{\mu}_{R}\|_{p},$
where the constants in $\asymp$ depend on $r,R\in(0,1)$.
Lemma 3.4.
[19, Lemma 5.1] Suppose $r\in(0,1)$ and $0<p<\infty$. Then we have
$\int_{\Omega}|f(z)|^{p}d\mu(z)\lesssim\int_{\Omega}|f(z)|^{p}\widehat{\mu}_{r}(z)dV(z)$
for all $f\in{\mathcal{H}}(\Omega)$.
Now we give the proof of Theorem 1.2
Proof of Theorem 1.2.
Thanks to the Montel’s theorem and Lemma 3.3, we need to prove (ii) $\implies$ (iii) $\implies$ (iv) $\implies$ (ii), (ii) $\implies$ (i).
(ii) $\implies$ (iii). For $\{c_{k}\}\in l^{p}$, let
$f(z)=\sum_{k=0}^{\infty}c_{k}K_{\Omega}(z,w_{k})^{N}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}.$
By Lemma 3.1, then $f\in A_{\alpha}^{p}$ and $\|f\|_{A_{\alpha}^{p}}^{p}\lesssim\sum_{k=0}^{\infty}|c_{k}|^{p}$. Since $J_{\mu}$ is bounded, hence
(3.3) $\begin{aligned} &\int_{\Omega}\bigg|\sum_{k=0}^{\infty}c_{k}K_{\Omega}(z,w_{k})^{N}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}\bigg|^{q}d\mu(z)\\ &\leq\|\mu\|^{q}\|f\|_{A_{\alpha}^{p}}^{q}\\ &\lesssim\|\mu\|^{q}\left(\sum_{k=0}^{\infty}|c_{k}|^{p}\right)^{\frac{q}{p}}.\end{aligned}$
Let
$\begin{cases}r_{0}(t)=\begin{cases}1,\,\,0\leq t-[t]\frac{1}{2},\\ -1,\,\,\frac{1}{2}\leq t-[t]<1,\end{cases}\\ r_{k}(t)=r_{0}(2^{k}t),\,\,k=1,2,3,\dots\end{cases}$
be Rademacher functions on $(0,1)$ and set
$f_{t}(z)=\sum_{k=0}^{\infty}r_{k}(t)c_{k}K_{\Omega}(z,w_{k})^{N}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}.$
Replacing $f(z)$ by $f_{t}(z)$ in (3.3), integrate with respect to $t$ from 0 to $1$, by Fubini’s theorem and Khinchine’s inequality, we obtain that
(3.4) $\begin{aligned} &\int_{\Omega}\left(\sum_{k=0}^{\infty}|c_{k}|^{2}K_{\Omega}(w_{k},w_{k})^{2(\frac{1-\alpha}{p}-N)}|K_{\Omega}(z,w_{k})|^{2N}\right)^{\frac{q}{2}}d\mu(z)\\ &\lesssim\int_{\Omega}\int_{0}^{1}\bigg|\sum_{k=0}^{\infty}r_{k}(t)c_{k}K_{\Omega}(z,w_{k})^{N}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}\bigg|^{q}dtd\mu(z)\\ &=\int_{0}^{1}\int_{\Omega}\bigg|\sum_{k=0}^{\infty}r_{k}(t)c_{k}K_{\Omega}(z,w_{k})^{N}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}\bigg|^{q}d\mu(z)dt\\ &\lesssim\|\mu\|^{q}\left(\sum_{k=0}^{\infty}|c_{k}|^{p}\right)^{\frac{q}{p}}.\end{aligned}$
On the one hand, if $q\geq 2$, then
$\int_{\Omega}\left(\sum_{k=0}^{\infty}|c_{k}|^{2}K_{\Omega}(w_{k},w_{k})^{2(\frac{1-\alpha}{p}-N)}|K_{\Omega}(z,w_{k})|^{2N}\right)^{\frac{q}{2}}d\mu(z)$
$\geq\int_{\Omega}\left(\sum_{k=0}^{\infty}|c_{k}|^{2}K_{\Omega}(w_{k},w_{k})^{2(\frac{1-\alpha}{p}-N)}|K_{\Omega}(z,w_{k})|^{2N}\chi_{k}(z)\right)^{\frac{q}{2}}d\mu(z)$
$\geq\int_{D_{k}}\sum_{k=0}^{\infty}|c_{k}|^{q}K_{\Omega}(w_{k},w_{k})^{q(\frac{1-\alpha}{p}-N)}|K_{\Omega}(z,w_{k})|^{qN}d\mu(z).$
On the other hand, if $0<q<2$, then Hölder’s inequality yields that
$\int_{D_{k}}\sum_{k=0}^{\infty}|c_{k}|^{q}K_{\Omega}(w_{k},w_{k})^{q(\frac{1-\alpha}{p}-N)}|K_{\Omega}(z,w_{k})|^{qN}d\mu(z)$
$=\int_{\Omega}\sum_{k=0}^{\infty}|c_{k}|^{q}K_{\Omega}(w_{k},w_{k})^{q(\frac{1-\alpha}{p}-N)}|K_{\Omega}(z,w_{k})|^{qN}\chi_{k}(z)d\mu(z)$
$\leq\int_{\Omega}\left(\sum_{k=0}^{\infty}|c_{k}|^{2}K_{\Omega}(w_{k},w_{k})^{2(\frac{1-\alpha}{p}-N)}|K_{\Omega}(z,w_{k})|^{2N}\right)^{\frac{q}{2}}\left(\sum_{k=0}^{\infty}\chi_{k}(z)\right)^{1-\frac{q}{2}}d\mu(z)$
$\leq N^{1-\frac{q}{2}}\int_{\Omega}\left(\sum_{k=0}^{\infty}|c_{k}|^{2}K_{\Omega}(w_{k},w_{k})^{2(\frac{1-\alpha}{p}-N)}|K_{\Omega}(z,w_{k})|^{2N}\right)^{\frac{q}{2}}d\mu(z).$
In a word
$\int_{D_{k}}\sum_{k=0}^{\infty}|c_{k}|^{q}K_{\Omega}(w_{k},w_{k})^{q(\frac{1-\alpha}{p}-N)}|K_{\Omega}(z,w_{k})|^{qN}d\mu(z)$
$\lesssim\left(\sum_{k=0}^{\infty}|c_{k}|^{2}K_{\Omega}(w_{k},w_{k})^{2(\frac{1-\alpha}{p}-N)}|K_{\Omega}(z,w_{k})|^{2N}\right)^{\frac{q}{2}}d\mu(z)$
holds for all $0<q<\infty$. By (2.2) and (3.4), we obtain that
(3.5) $\sum_{k=0}^{\infty}|c_{k}|^{q}K_{\Omega}(w_{k},w_{k})^{q(\frac{1-\alpha}{p})}\mu(D_{k})\lesssim\|\mu\|^{q}\left(\sum_{k=0}^{\infty}|c_{k}|^{p}\right)^{\frac{q}{p}}.$
Let $a_{k}=|c_{k}|^{q}$, then $\{a_{k}\}\in l^{\frac{p}{q}}$. The fact that $(l^{\frac{p}{q}})^{*}=l^{\frac{p}{p-q}}$, Lemma 2.6 and (3.5) yield
$\|\mu\|$ $\gtrsim\|\mu(D_{k})K_{\Omega}(w_{k},w_{k})^{q(\frac{1-\alpha}{p})}\|_{l^{\frac{p}{p-q}}}^{\frac{1}{q}}$
$\asymp\|\widehat{\mu}_{r}(B_{\Omega}(w_{k},r))p_{\Omega}(w_{k})^{2+\frac{2q(\alpha-1)}{p}}\|_{l^{\frac{p}{p-q}}}^{\frac{1}{q}},$
which give the proof of (ii) $\implies$ (iii) if replace $w_{k}$ by some $r$-lattice $\{a_{j}\}$.
[…]
Let $\{a_{j}\}_{j=1}^{\infty}$ be some $r$-lattice such that $B_{\Omega}(z,r)\subset B_{\Omega}(a_{j},r)$ for all $z\in B_{\Omega}(a_{j},r)$. For all $z\in B_{\Omega}(a_{j},r)$, Lemma 2.2 and (2.1) yield that
(3.6) $\begin{aligned} \widehat{\mu}_{R}(z)&\approx p_{\Omega}(z)^{-2}\mu\left(B_{\Omega}(z,R)\right)\\ &\lesssim p_{\Omega}(a_{j})^{-2}\mu\left(B_{\Omega}(a_{j},r)\right)\approx\widehat{\mu}_{r}(a_{j})\end{aligned}$
Lemma 2.2, 2.4 and (3.6) yield that
$\left\|p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widehat{\mu}_{R}(z)\right\|_{p-q}^{\frac{p}{p-q}}$ $\lesssim\sum_{j=1}^{\infty}\int_{B_{\Omega}(a_{j},r)}p_{\Omega}(z)^{\frac{2\alpha q}{p-q}}\widehat{\mu}_{R}(z)^{\frac{p}{p-q}}\,dV(z)$
$\lesssim\sum_{j=1}^{\infty}p_{\Omega}(a_{j})^{2+\frac{2\alpha q}{p-q}}\widehat{\mu}_{r}(a_{j})^{\frac{p}{p-q}}$
$=\left\|\left\{p_{\Omega}(a_{j})^{2+\frac{2(\alpha-1)q}{p}}\widehat{\mu}_{r}(a_{j})\right\}_{j=1}^{\infty}\right\|_{l^{\frac{p}{p-q}}}^{\frac{p}{p-q}}<\infty.$
(iv) $\implies$ (ii). Since $p>q$, hence $\frac{1}{\frac{p}{q}}+\frac{1}{\frac{p}{p-q}}=1$. For $f\in A_{\alpha}^{p}$, Lemma 3.4 and Lemma 2.6 yield that
$\int_{\Omega}|f(z)|^{q}d\mu(z)$
$\lesssim\int_{\Omega}|f(z)|^{q}\widehat{\mu}_{R}(z)dV(z)$
$\leq\bigg\||f(z)|^{q}K_{\Omega}(z,z)^{\frac{\alpha q}{p}}\Big\|_{\frac{p}{q}}\cdot\left\|K_{\Omega}(z,z)^{-\frac{\alpha q}{p}}\widehat{\mu}_{R}(z)\right\|_{\frac{p}{p-q}}$
$\asymp\left\|p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widehat{\mu}_{R}(z)\right\|_{\frac{p}{p-q}}\|f\|_{A_{\alpha}^{p}}^{q}$
$\lesssim\|f\|_{A_{\alpha}^{p}}^{q}.$
(ii) $\implies$ (i). Let $\{f_{j}\}$ be any bounded sequence in $A_{\alpha}^{p}$ and $f_{j}\to 0$ uniformly on each compact subset of $\Omega$ as $j\to\infty$. Since $p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widehat{\mu}_{R}(z)\in L^{\frac{p}{p-q}}$, hence there exists a set $K_{R}\subset\Omega$ such that
(3.7) $\int_{\Omega\setminus K_{R}}\bigg|p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widehat{\mu}_{R}(z)\bigg|^{\frac{p}{p-q}}dV(z)<\varepsilon.$
Consider
$\int_{\Omega}|f_{j}(z)|^{q}d\mu(z)$ $=\left(\int_{K_{R}}+\int_{\Omega\setminus K_{R}}\right)|f_{j}(z)|^{q}d\mu(z)$
$=I_{1}+I_{2}.$
For $I_{1}$, since $f_{j}\to 0$ uniformly on each compact subset of $\Omega$, then we have
$I_{1}=\int_{K_{R}}|f_{j}(z)|^{q}\leq C\sup_{z\in K_{R}}|f_{j}(z)|^{q}<\frac{\varepsilon}{2},\quad j\to\infty.$
Let $\mu_{R}$ be the restriction of $\mu$ to $\Omega\setminus K_{R}$, noting that $J_{\mu}$ is bounded if and only if $p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widehat{\mu}_{R}\in L^{\frac{p}{p-q}}$, by (3.7) we obtain that
$I_{2}=$ $\int_{\Omega\setminus K_{R}}|f_{j}(z)|^{q}d\mu(z)$
$=\int_{\Omega}|f_{j}(z)|^{q}d\mu(z)$
$\lesssim\left\|p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widehat{\mu}_{R}(z)\right\|_{\frac{p}{p-q}}\|f\|_{A_{\alpha}^{p}}^{q}$
$\lesssim\frac{\varepsilon}{2}.$
Therefore,
$\lim_{j\to\infty}\int_{\Omega}|f_{j}(z)|^{q}d\mu(z)=0,$
which means that $\mu$ is a vanishing $q$-Carleson measure, hence $J_{\mu}$ is compact. The proof is completed. ∎
Acknowledgments
This research work was supported by the National Natural Science Foundation of China (Grant No. 12261023, 11861023.)

### v2

@@
-A note for Carleson measure on bounded $\mathbb{C}$}-doconvex domains
+A note for Carleson measure on bounded $\mathbb{C}$}-convex domains
 Abstract.
 Let $0<q<p<\infty$, $\Omega$ be a bounded $\mathbb{C}$}-convex domains in ${\mathbb{C}}^{n}$.
-We establish several equivalent characterizations for the boundedness of Carleson embedding $J_{\mu}:A_{\alpha}^{p}\hookrightarrow L^{q}(\mu)$ on $\Omega$ with sharp $\mathcal{B}$} condition.
+We establish several equivalent characterizations for the boundedness of Carleson embedding $J_{\mu}:A_{\alpha}^{p}\hookrightarrow L^{q}(\mu)$ on $\Omega$ with sharp $\mathcal{B}$} type condition.
 Furthermore, we prove that the boundedness of $J_{\mu}$ is equivalent to its compactness.
@@
 Primary 46E30; Secondary 32A25, 32A36
-*:Corresponding author.
+*Corresponding author.
 1. Introduction and main result
@@
 The extension of $q$-Carleson measure to Bergman spaces has attracted considerable attention from many scholars, it was characterized by Hastings [18] and Oleinik [19] for $p\leq q$, Luecking [16] for $p>q$.
-In higher dimension, Cima-Wogen [7] and Luecking-Duren-Weir [14, 10] stuied the case of unit ball ${\mathbb{B}}^{n}$ by using the Euclidean geometry of ${\mathbb{B}}^{n}$ and conformal invariants (such as Bergman or Kobayashi balls) repectively.
-In order to extend the above results to other domains of ${\mathbb{C}}^{n}$, Cima-Mercer and Abate-Raissy-Saracco characterrized the $q$-Carleson measure for $A^{p}$ on strongly pseudoconvex domains by using the Euclidean geometric quantities and intrinsic Kobayashi geometry respectively, see [1, 2, 3, 8] for the details.
+In higher dimension, Cima-Wogen [7] and Luecking-Duren-Weir [14, 10] stuied the case of unit ball ${\mathbb{B}}^{n}$ by using the Euclidean geometry of ${\mathbb{B}}^{n}$ and conformal invariants (such as Bergman or Kobayashi balls) respectively.
+In order to extend the above results to other domains of ${\mathbb{C}}^{n}$, Cima-Mercer and Abate-Raissy-Saracco characterized the $q$-Carleson measure for $A^{p}$ on strongly pseudoconvex domains by using the Euclidean geometric quantities and intrinsic Kobayashi geometry respectively, see [1, 2, 3, 8] for the details.
 For the case of weakly pseudoconvex domains, Li-Liu-Wang [13] characterized the $2$-Carleson measurable for $A^{2}$ by using Kobayashi balls on bounded convex domains with smooth boundary of finite type.
 Bi-Su-Zhang [5] characterized the $2$-Carleson measure for $A^{2}$ and $A^{p}$ respectively on bounded $\mathbb{C}$}-convex domains.
-Khanh-Liu-Thuc [11] introduced the sharp $\mathcal{B}$} type condition (see Sect.2.4 ) for Bergman kerenl.
-Roughly speaking, Bergman kernel satisfies the sharp $\mathcal{B}$} condition ensures that the Bergman projection is bounded on $L^{p}$ for $1<p<\infty$.
+Khanh-Liu-Thuc [11] introduced the sharp $\mathcal{B}$} type condition (see Sect.2.4 below) for Bergman kernel.
+Roughly speaking, Bergman kernel satisfies the sharp $\mathcal{B}$} type condition ensures that the Bergman projection is bounded on $L^{p}$ for $1<p<\infty$.
 More recently, there is a series of work involved function spaces and operator theory related such domains.
-For example, Khanh-Liu-Thuc-Tien [11, 12] characterrized the boundedness and Schatten class of Bergman-Toeplitz operators for weighted spaces on weakly pseudoconvex domains with Bergman kernel satisfies sharp $\mathcal{B}$} condition.
-Zhang [19] characterized the $q$-Carleson measure for the weighted Bergman spaces $A_{\alpha}^{p}$ on bounded $\mathbb{C}$}-convex domains with sharp $\mathcal{B}$} type condition, and what’s more, Zhang’s results extended the results in [5] to the case of $p\leq q$ and the weighted Bergman spaces $A_{\alpha}^{p}$, but for the case of $p>q$, Zhang didn’t characterized the full Carleson measure, Zhang’s result is stated as follows.
+For example, Khanh-Liu-Thuc-Tien [11, 12] characterized the boundedness and Schatten class of Bergman-Toeplitz operators for weighted spaces on weakly pseudoconvex domains with Bergman kernel satisfies sharp $\mathcal{B}$} type condition.
+Zhang [20] characterized the $q$-Carleson measure for the weighted Bergman spaces $A_{\alpha}^{p}$ on bounded $\mathbb{C}$}-convex domains with sharp $\mathcal{B}$} type condition, and what’s more, Zhang’s results extended the results in [5] to the case of $p\leq q$ and the weighted Bergman spaces $A_{\alpha}^{p}$, but for the case of $p>q$, Zhang didn’t characterize the full Carleson measure, Zhang’s result is stated as follows.
 Theorem 1.1.
-[19, Theorem 1.8] Let $\Omega\subset{\mathbb{C}}^{n}$ be a bounded $\mathbb{C}$}-convex domain whose Bergman kernel of sharp $\mathcal{B}$} type, $\mu$ be a finite positive Borel measure on $\Omega$.
+[20, Theorem 1.8] Let $\Omega\subset{\mathbb{C}}^{n}$ be a bounded $\mathbb{C}$}-convex domain whose Bergman kernel satisfies sharp $\mathcal{B}$} type condition, $\mu$ be a finite positive Borel measure on $\Omega$.
 Assume that $0<q<p<\infty$, if $\mu$ satisfies one of the following three conditions
@@
 Theorem 1.2.
-Let $\Omega$ be a bounded $\mathbb{C}$}-convex domain whose Bergman kernel of sharp $\mathcal{B}$}-type.
+Let $\Omega\subset{\mathbb{C}}^{n}$ be a bounded $\mathbb{C}$}-convex domain whose Bergman kernel of sharp $\mathcal{B}$} type.
 Let $\mu$ be a finite positive Borel measure, $0<q<p<\infty$.
-Then the following are equivalent.
+Then the following statements are equivalent.
   • (i)
@@
 Given a domain $\Omega\subset\mathbb{C}^{n}$ (with $n\geq 2$), the (infinitesimal) Kobayashi metric is the pseudo-Finsler metric defined by
-$\kappa_{\Omega}(x;v):=\inf\bigl\{|\xi|:f\in{\mathcal{H}}(\mathbb{D},\Omega),\text{ with }f(0)=x,\;(df)_{0}(\xi)=v\bigr\}.$
-For any curve $\sigma:[a,b]\to\Omega$, the Kobayashi length is
+$\kappa_{\Omega}(x;v):=\inf\bigl\{|\xi|:f\in H(\mathbb{D},\Omega),\text{ with }f(0)=x,\;(df)_{0}(\xi)=v\bigr\},$
+where $H({\mathbb{D}},\Omega)$ denotes the holomorphic section from unit disc $\mathbb{D}$} to $\Omega$. For any curve $\sigma:[a,b]\to\Omega$, the Kobayashi length is
 $l_{\kappa}(\sigma):=\int_{a}^{b}\kappa_{\Omega}\bigl(\sigma(t);\sigma^{\prime}(t)\bigr)\,dt,$
@@
 $\sigma(a)=x\text{ and }\sigma(b)=y\bigr\}.$
-Suppose $\Omega$ is a bounded domains in ${\mathbb{C}}^{n}$, $z\in\Omega$ and $r\in(0,1)$. Let $B_{\Omega}(z,r)$ be the Kobayashi balls of centre $z$ and radius $\frac{1}{2}\log\frac{1+r}{1-r}$, i.e.
+Suppose $\Omega$ is a bounded domain in ${\mathbb{C}}^{n}$, $z\in\Omega$ and $r\in(0,1)$. Let $B_{\Omega}(z,r)$ be the Kobayashi balls of centre $z$ and radius $\frac{1}{2}\log\frac{1+r}{1-r}$, i.e.
 $B_{\Omega}(z,r)=\bigl\{z\in X\mid\tanh d_{\Omega}^{K}(z,w)<r\bigr\}.$
 Note that $\rho_{\Omega}=\tanh d_{\Omega}^{K}$ is still a distance on $X$ since $\tanh$ is a strictly convex function on $\mathbb{R}^{+}$.
-For a bounded pseudoconvex domains $\Omega\subset{\mathbb{C}}^{n}$, Set $A^{2}$ be the Hilbert spaces consists of all $f\in{\mathcal{H}}(\Omega)$ such that $\|f\|_{A^{2}}^{2}=\int_{\Omega}|f(z)|^{2}dV(z)<\infty$, where $dV(z)$ is the ordinary Lebesgue measure.
+For a bounded pseudoconvex domain $\Omega\subset{\mathbb{C}}^{n}$, Set $A^{2}$ be the Hilbert spaces consists of all $f\in{\mathcal{H}}(\Omega)$ such that $\|f\|_{A^{2}}^{2}=\int_{\Omega}|f(z)|^{2}dV(z)<\infty$, where $dV(z)$ is the ordinary Lebesgue measure.
 Riesz’s theorem yields that there exists a unique reproducing kernel function $K_{\Omega}(z,w)$ such that
@@
 $H_{2}:=q+\operatorname{Span}(q_{1}-q,q_{2}-q)^{\perp},\quad\Omega_{2}:=\Omega\cap H_{2},$
-and. Continuing this procedure we are led to an orthonormal basis
+and continuing this procedure we are led to an orthonormal basis
 $e_{j}=\frac{q_{j}-q}{|q_{j}-q|},\quad 1\leq j\leq n,$
@@
 Lemma 2.2.
-[NIkolv2011, Theorem 1] Let $\Omega\subset\mathbb{C}^{n}$ be a $\mathbb{C}$-convex domain containing no complex lines, and $z\in\Omega$. Then for any $r\in(0,1)$,
+[17, Theorem 1] Let $\Omega\subset\mathbb{C}^{n}$ be a $\mathbb{C}$-convex domain containing no complex lines, and $z\in\Omega$. Then for any $r\in(0,1)$,
 $\mathbb{D}^{n}(z,c_{1}\eta(z))\subset B_{\Omega}(z,r)\subset\mathbb{D}^{n}(z,c_{2}\eta(z)),$
@@
 [5, Proposition 2.3] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded $\mathbb{C}$-convex domain.
-For any $r\in(0,1)$, we denote $R:=\frac{1}{2}(1+r)\in(0,1)$.
+For any $r\in(0,1)$, we denote $R:=\frac{1}{2}(1+r)$.
 Then for any $z_{0}\in\Omega$ and $z\in B_{\Omega}(z_{0},r)$, we have
@@
 for all nonnegative plurisubharmonic functions $\varphi$ on $\Omega$.
-The following cover type lemma plays an significant role in the proof of Theorem 1.2.
+The following cover type lemma plays a significant role in the proof of Theorem 1.2.
 Lemma 2.4.
@@
 Lemma 2.5.
-[19, Lemma 2.4] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded $\mathbb{C}$-convex domain. Then for any $r\in(0,1)$, $z\in\Omega$ and $w\in B_{\Omega}(z,r)$, we have that
+[20, Lemma 2.4] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded $\mathbb{C}$-convex domain. Then for any $r\in(0,1)$, $z\in\Omega$ and $w\in B_{\Omega}(z,r)$, we have
 $|K_{\Omega}(z,w)|\gtrsim K_{\Omega}(z,z)^{\frac{1}{2}}K_{\Omega}(w,w)^{\frac{1}{2}},$
 where the constant in $\gtrsim$ depends only on $r$ and the domain $\Omega$.
-For $z\in\Omega$ and $r\in(0,1)$. Let $p_{\Omega}(z)=\prod\limits_{i=1}^{n}\eta_{i}(z)$, then by [19], we have
-(2.1) $\frac{1}{C}p_{\Omega}(z)\leq p_{\Omega}(w)\leq Cp_{\Omega}(z),\quad w\in B_{\Omega}(z,r)$
-for some positive constant $C=C(r)$.
+For $z\in\Omega$ and $r\in(0,1)$. Let $p_{\Omega}(z)=\prod\limits_{i=1}^{n}\eta_{i}(z)$, then by [20], for some positive constant $C=C(r)$ we have
+(2.1) $\frac{1}{C}p_{\Omega}(z)\leq p_{\Omega}(w)\leq Cp_{\Omega}(z),\quad w\in B_{\Omega}(z,r).$
 By (2.1), the following estimate for Bergman kernel on diagonal is given.
 Lemma 2.6.
-[19, Lemma 2.5] Let $\Omega\subset\mathbb{C}^{n}$ be a $\mathbb{C}$-convex domain containing no complex lines. Then
+[20, Lemma 2.5] Let $\Omega\subset\mathbb{C}^{n}$ be a $\mathbb{C}$-convex domain containing no complex lines. Then
 $K_{\Omega}(z,z)\asymp\rho_{\Omega}(z)^{-2},\quad z\in\Omega.$
@@
 (2.2) $|K_{\Omega}(z,w)|\gtrsim K_{\Omega}(z,z).$
-2.4. Bergman kernel with sharp $\mathcal{B}$} conditions
-In this subsection, we introduce a sharp $\mathcal{B}$} condition for Bergman kerenl, we firstly recall that the definition of $\mathcal{B}$}- systems.
+2.4. Bergman kernel with sharp $\mathcal{B}$} type conditions
+In this subsection, we introduce a sharp $\mathcal{B}$} type condition for Bergman kernel, we firstly recall that the definition of $\mathcal{B}$}- systems.
 Let $\Omega$ be a bounded domain in $\mathbb{C}^{n}$.
-For $z\in\overline{\Omega}$ near the boundary $\partial\Omega$, a family of functions $\mathcal{B}=\{b_{j}(z,\cdot)\}_{j=1}^{n}$ is called a $\mathcal{B}$-system at $z$ if there exist a neighbourhood $U$ of $z$ and a positive integer $m\geq 2$ such that for all $w\in U$,
-$b_{1}(z,w):=\frac{1}{\delta(z,w)}\quad\text{and}\quad b_{j}(z,w):=\sum_{k=2}^{m}\left(\frac{A_{jk}(z)}{\delta(z,w)}\right)^{\frac{1}{k}},\quad\text{for }j=2,\ldots,n,$
+For $z\in\overline{\Omega}$ near the boundary $\partial\Omega$, a family of functions $\mathcal{B}=\{b_{j}(z,\cdot)\}_{j=1}^{n}$ is called a $\mathcal{B}$-system at $z$ if there exist a neighborhood $U$ of $z$ and a positive integer $m\geq 2$ such that for all $w\in U$,
+$b_{1}(z,w):=\frac{1}{\delta(z,w)}\quad\text{and}\quad b_{j}(z,w):=\sum_{k=2}^{m}\left(\frac{A_{jk}(z)}{\delta(z,w)}\right)^{\frac{1}{k}},\quad j=2,\ldots,n,$
 where $A_{jk}$ are non-negative bounded functions, and $\delta(z,w)$ is the pseudo-distance between $z$ and $w$ defined as follows:
-$\delta(z,w):=|\delta_{\Omega}(z)|+|\delta_{\Omega}(w)|+|z_{1}-w_{1}|+\sum_{l=2}^{n}\sum_{s=2}^{m}A_{ls}(z)|z_{l}-w_{l}|^{s}$
+$\delta(z,w):=|\delta_{\Omega}(z)|+|\delta_{\Omega}(w)|+|z_{1}-w_{1}|+\sum_{l=2}^{n}\sum_{s=2}^{m}A_{ls}(z)|z_{l}-w_{l}|^{s}.$
 For $z\in\overline{\Omega}$ near the boundary, the Bergman kernel $K_{\Omega}$ is said to be of $\mathcal{B}$-type at $z\in\overline{\Omega}$ near the boundary $\partial\Omega$ if there exist positive constants $c$ and $C$ such that for all $w\in\Omega\cap\mathbb{B}(z,c)$,
 $|K_{\Omega}(z,w)|\leq C\prod_{j=1}^{n}b_{j}^{2}(z,w).$
-The Bergman kernel $K_{\Omega}$ is said to be of sharp $\mathcal{B}$-type at $z$ if $K_{\Omega}$ is of $\mathcal{B}$-type and has the sharp lower-bound on diagonal, i.e.
+The Bergman kernel $K_{\Omega}$ is said to be of sharp $\mathcal{B}$ type at $z$ if $K_{\Omega}$ is of $\mathcal{B}$-type and has the sharp lower-bound on diagonal, i.e.
 $K_{\Omega}(z,z)\asymp\prod_{j=1}^{n}b_{j}^{2}(z,z).$
-We finish this section with a Forelli-Rudin estimate on domains with Bergman kernel satisfying the sharp $\mathcal{B}$} condition, which will be used to construct an atomic type decomposition.
+We finish this section with a Forelli-Rudin estimate on domains with Bergman kernel satisfying the sharp $\mathcal{B}$} type condition, which will be used to construct an atomic type decomposition.
 Lemma 2.7.
-[19, Proposition 2.7] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain whose Bergman kernel $K_{\Omega}$ is of sharp $\mathcal{B}$-type.
+[20, Proposition 2.7] Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain whose Bergman kernel $K_{\Omega}$ is of sharp $\mathcal{B}$ type.
 Assume that $\alpha\leq 0,\,\beta\geq 0,\,t+\alpha\geq 1$ and $1<\beta-2\alpha<2t-2$.
@@
 For $0<p<\infty$ and $\{c_{k}\}\in l^{p}$, $\alpha+Np>1$.
-Set $f(z)=\sum_{k=1}^{\infty}c_{k}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}K_{\Omega}(z,w_{k})^{N}$,
-$\|f\|_{A_{\alpha}^{p}}\lesssim 1.$
+Set $f(z)=\sum\limits_{k=1}^{\infty}c_{k}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}K_{\Omega}(z,w_{k})^{N}$, then we have
+$\|f\|_{A_{\alpha}^{p}}\lesssim\|c_{k}\|_{l^{p}}\asymp 1.$
 The following lemma can be obtained by Lemma 2.7 and Schur’s test, it can be used to construction of atomic decomposition in the proof of Theorem 3.1.
 Lemma 3.2.
-Let $\Omega$ be a bounded pseudoconvex domain whose Bergman kernel $K_{\Omega}(z,w)$ satisfying sharp $\mathcal{B}$} condition. Assume that
+Let $\Omega\subset{\mathbb{C}}^{n}$ be a bounded pseudoconvex domain whose Bergman kernel $K_{\Omega}(z,w)$ satisfies sharp $\mathcal{B}$} type condition. Assume that
 (3.1) $\begin{cases}&q+p\sigma+\alpha\leq 0,\\ &s+q+p\sigma+\alpha\geq 1,\\ &-1<-2(q+p\sigma+\alpha)<2s-2,\\ &t+p^{\prime}\sigma\leq 0,\\ &s+t+p^{\prime}\sigma\geq 1,\\ &-1<-2(t+p^{\prime}\sigma+\alpha)<2s-2,\\ &q+s+t-1\leq 0.\end{cases}$
@@
 Proof.
-To prove the boundedness of $T_{q,s,t}$, we only prove the boundedness $S_{q,s,t}f(z)=K_{\Omega}(z,z)^{q}\int_{\Omega}|K_{\Omega}(z,w)|^{s}K_{\Omega}(w,w)^{t}f(w)dV(w)$.
+To prove the boundedness of $T_{q,s,t}$, we only prove the boundedness of $S_{q,s,t}f(z)=K_{\Omega}(z,z)^{q}\int_{\Omega}|K_{\Omega}(z,w)|^{s}K_{\Omega}(w,w)^{t}f(w)dV(w)$.
 To do this, let $h(z)=K_{\Omega}(z,z)^{\sigma}$,
@@
 $S_{q,s,t}F(z)\gtrsim\sum_{k=0}^{\infty}|c_{k}|K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}\gtrsim|f(z)|.$
-Hence, lemma 3.1 that yields
+Hence, lemma 3.2 yields that
 $\|f\|_{\alpha}^{p}\lesssim\|S_{q,s,t}F\|_{L_{\alpha}^{p}}^{p}\lesssim\|F\|_{L_{\alpha}^{p}}^{p}\lesssim\|c_{k}\|_{l^{p}}^{p}.$
 The proof is completed. ∎
-In order to prove (iv) $\Longleftrightarrow$ (v) in Theorem LABEL:restate, we need the following results.
+In order to prove (iv) $\Longleftrightarrow$ (v) in Theorem 1.2, we need the following results.
 Lemma 3.3.
-[19, Lemma 5.3] Let $1<p<\infty$ and $s\in\mathbb{R}$.
+[20, Lemma 5.3] Let $1<p<\infty$ and $s\in\mathbb{R}$.
 Then the function $p_{\Omega}^{s}\widehat{\mu}_{r}\in L^{p}(\Omega)$ for some $r\in(0,1)$ if and only if $p_{\Omega}^{s}\widehat{\mu}_{R}\in L^{p}(\Omega)$ for all $R\in(0,1)$.
@@
 Lemma 3.4.
-[19, Lemma 5.1] Suppose $r\in(0,1)$ and $0<p<\infty$. Then we have
+[20, Lemma 5.1] Suppose $r\in(0,1)$ and $0<p<\infty$. Then we have
 $\int_{\Omega}|f(z)|^{p}d\mu(z)\lesssim\int_{\Omega}|f(z)|^{p}\widehat{\mu}_{r}(z)dV(z)$
@@
 Proof of Theorem 1.2.
-Thanks to the Montel’s theorem and Lemma 3.3, we need to prove (ii) $\implies$ (iii) $\implies$ (iv) $\implies$ (ii), (ii) $\implies$ (i).
+Thanks to the Montel’s theorem and Lemma 3.3, we only need to prove (ii) $\implies$ (iii) $\implies$ (iv) $\implies$ (ii), (ii) $\implies$ (i).
 (ii) $\implies$ (iii). For $\{c_{k}\}\in l^{p}$, let
 $f(z)=\sum_{k=0}^{\infty}c_{k}K_{\Omega}(z,w_{k})^{N}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}.$
-By Lemma 3.1, then $f\in A_{\alpha}^{p}$ and $\|f\|_{A_{\alpha}^{p}}^{p}\lesssim\sum_{k=0}^{\infty}|c_{k}|^{p}$. Since $J_{\mu}$ is bounded, hence
+By theorem 3.1, then $f\in A_{\alpha}^{p}$ and $\|f\|_{A_{\alpha}^{p}}^{p}\lesssim\sum_{k=0}^{\infty}|c_{k}|^{p}$. Since $J_{\mu}$ is bounded, hence
 (3.3) $\begin{aligned} &\int_{\Omega}\bigg|\sum_{k=0}^{\infty}c_{k}K_{\Omega}(z,w_{k})^{N}K_{\Omega}(w_{k},w_{k})^{\frac{1-\alpha}{p}-N}\bigg|^{q}d\mu(z)\\ &\leq\|\mu\|^{q}\|f\|_{A_{\alpha}^{p}}^{q}\\ &\lesssim\|\mu\|^{q}\left(\sum_{k=0}^{\infty}|c_{k}|^{p}\right)^{\frac{q}{p}}.\end{aligned}$
@@
 For all $z\in B_{\Omega}(a_{j},r)$, Lemma 2.2 and (2.1) yield that
-(3.6) $\begin{aligned} \widehat{\mu}_{R}(z)&\approx p_{\Omega}(z)^{-2}\mu\left(B_{\Omega}(z,R)\right)\\ &\lesssim p_{\Omega}(a_{j})^{-2}\mu\left(B_{\Omega}(a_{j},r)\right)\approx\widehat{\mu}_{r}(a_{j})\end{aligned}$
-Lemma 2.2, 2.4 and (3.6) yield that
+(3.6) $\begin{aligned} \widehat{\mu}_{R}(z)&\approx p_{\Omega}(z)^{-2}\mu\left(B_{\Omega}(z,R)\right)\\ &\lesssim p_{\Omega}(a_{j})^{-2}\mu\left(B_{\Omega}(a_{j},r)\right)\approx\widehat{\mu}_{r}(a_{j}).\end{aligned}$
+Lemma 2.2, lemma 2.4 and (3.6) yield that
 $\left\|p_{\Omega}(z)^{\frac{2\alpha q}{p}}\widehat{\mu}_{R}(z)\right\|_{p-q}^{\frac{p}{p-q}}$ $\lesssim\sum_{j=1}^{\infty}\int_{B_{\Omega}(a_{j},r)}p_{\Omega}(z)^{\frac{2\alpha q}{p-q}}\widehat{\mu}_{R}(z)^{\frac{p}{p-q}}\,dV(z)$